OpenStudy (anonymous):

Can anyone explain how this: p(z) = 2 + 5z - 3z^2 factors into this: -3(z + 1/3)(z-2) ? This seems to be a similar problem to the following: 7x^2 -5x -2 becomes: 7(x - 1)(x + 2/7) I don't understand the steps required to reach these solutions. :(

OpenStudy (mayankdevnani):

by applying factor theorm..

OpenStudy (mayankdevnani):

@math0101

OpenStudy (anonymous):

-3x^2 + 5x + 2 factor -1 first

OpenStudy (mayankdevnani):

first of all , p(z) converts into 0 i.e p(z)=0[remainder theorem]

OpenStudy (mayankdevnani):

so, p(-2)=2+5(-2)-3(2)^2 p(-2)=12-12=0

OpenStudy (mayankdevnani):

therefore , -2 is a zero of a polynomial

OpenStudy (mayankdevnani):

and facrtorize it

OpenStudy (anonymous):

nope, still not getting it. :) Here's what I've got so far: \[2+5z-3z^2\] \[-3z^2+5z+2\] \[3(z^2+5z/3+2/3)\] \[-(5z/3)/2 +- \sqrt{(5z/3)/2-2/3}\] \[x1 = \sqrt{2/3}\] Well... somewhere here everything breaks down. :/

OpenStudy (anonymous):

Hmm, I think I'll rephrase the question.